OpenStudy (anonymous):

Solve the system of equations: what is the x-value for the solution? 2x + 7y = 3 and x = 1 - 4y

OpenStudy (anonymous):

To get the x-value, reverse solve for x here: x = 1 - 4y 4y + x = 1 Now, set up the system of equations: 2x + 7y = 3 x + 4y = 1 -4(2x + 7y = 3) 7(x + 4y = 1) -8x - 28y = -12 7x + 28y = 7 -x = -5 x = 5

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

np :)